# ProGeny II: the impact of libraries and model configurations on inferred galaxy properties in SED fitting

### v1

@@ 4.5 Impact on Derived Cosmic Star Formation History
We can use the volume-limited nature of our GAMA sample to compute the CSFH using the individual SFHs in the same manner as outlined by Carnall et al. (2019); Leja et al. (2019); Bellstedt et al. (2020b); Bellstedt et al. (2024), as a manner of assessing the systematic bias of the SFHs. The CSFH is well understood from high-$z$ observations from a number of works (including Madau & Dickinson, 2014; Driver et al., 2018; D’Silva et al., 2023), and therefore any deviation in the forensic101010‘Forensic’ here refers to the process of constructing a historical property (the CSFH) from present-day galaxies. reconstructed CSFH and the observed one points to underlying age biases in the SED fitting process.
@@
The top panel of Fig. 16 presents the derived CSFH using the fiducial implementation for all six SPLs analysed in this work, as compared with the compilation of measurements by Madau & Dickinson (2014) [^11] Measurements included in this compilation come from a suite of studies: Sanders (2003); Takeuchi et al. (2003); Wyder et al. (2005); Schiminovich et al. (2005); Dahlen et al. (2007); Reddy & Steidel (2009); Robotham & Driver (2011); Magnelli et al. (2011); Magnelli et al. (2013); Cucciati et al. (2012); Bouwens et al. (2012a, b); Schenker et al. (2013); Gruppioni et al. (2013) and the more recent measurement by D’Silva et al. (2023). The methodology of computing the CSFH from individual SFHs is identical to that outlined by Bellstedt et al. (2024), where the ProSpect-derived Vmax values are used to compute a volume correction for low-mass galaxies. The agreement between recovered CSFH curves using different SPLs is relatively good until 6 Gyr, beyond which the curves are more divergent, pointing to the underlying disagreements and uncertainties. This is consistent with the conclusions reached by Bravo et al. (2022), who determined that the recovery of galaxy properties from SED fitting was accurate to around 6 Gyr, beyond which modelling choices drove the recovered properties. It is notable that the CSFH derived using BC03, FSPS, and M05 all recover a substantial peak beyond 12 Gyr in lookback time, which is not reflected in the high-redshift of Madau & Dickinson (2014) and D’Silva et al. (2023) presented in the figure. Furthermore, the observed peak of the CSFH at a lookback time of $\sim$10 Gyr is not well recovered by BC03, M05, CB19 and ProGeny, instead being much flatter between 6-12 Gyr. The overall shape of the CSFH is best recovered using FSPS, however the overall normalisation is slightly low.
@@
The remaining panels of Fig. 16 present the derived CSFH using all tested SPLs with the alternative implementations explored in this work. From a visual comparison between the derived CSFH curves and the underlying observational measurements, it is clear that the fiducial implementation is better capable of recovering the observed CSFH. With the change to assuming a constant Z, it can be seen that the CSFH surves tend to start skewing toward younger peaks, which is a trend that continues further when moving to the $Z_{\odot}$ implementation in the third panel. It is also notable that there is a clear artefact recovered in the $Z_{\odot}$ implementation, where the SFRD values are overestimated around a Gyr in the past. The only SPL for which this behaviour is not present is M05 While the overall recovery from the b5 SFH configuration in the bottom panel is similar to the fiducial, the impact of the stepwise parametrisation is clear, in that the gradients are unable to be recovered. In the b5 SFH implementation it is noteworthy that the lowest-$z$ bin of the CSFH displays much more discrepancy between the SPL values than the other SED-fitting implementations. Note that the most recent burst of star formation available with the b5 SFH implementation hardly appears in this plot, as the CSFH only focuses on the region of lookback time that is covered by all galaxies, beyond $z=0.06$. This is an indication that the relative decoupling of the SFR values in neighbouring bins acts to reduce the constraint on slightly older stellar populations. In general, the quality of CSFH recovery by the b5 mode is similar to the recovery presented by Leja et al. (2019) using a stepwise parametric SFH with the code Prospector, where it was shown that the selection of a prior can also have a large impact on the SFH recovery.
@@
A numerical comparison of the overall CSFH recovery is shown in Fig. 17. The top panel presents the mean absolute deviation (MAD) from the D’Silva et al. (2023) CSFH for each SPL/configuration combination, computed over the lookback time range for which there are observations ($<$12.5 Gyr). The ProGeny fiducial run has the overall lowest MAD (only marginally better than FSPS). In general the fiducial configuration is best capable of replicating the CSFH, as indicated by the fact that the MAD averaged over SPLs (rightmost panel of Fig. 17) is the lowest for the fiducial implementation.
The second panel of Fig. 17 presents the offset of the SFRD at lookback times of 0. The grey shaded region indicates the range covered by the observational uncertainty of D’Silva et al. (2023). In general, the $Z_{\odot}$ implementation can be seen to produce too much star formation, whereas the fiducial, constant Z, and b5 SFH implementations tend to produce underestimates. Overall, M05, BC03 and ProGeny produce a present-day SFRD that is most consistent with observations. This is perhaps unsurprising, as BC03 was used to infer the observational results by D’Silva et al. (2023).
The third panel shows the offset of the SFRD at the observed peak of the CSFH. It is notable that the magnitude of the offsets at $\sim$10 Gyr in lookback time are substantially larger than present day values (by a factor of $\sim$6), demonstrating the increased difficulty of constraining star formation at older times. All SPLs underestimate the value of the peak, however FSPS is the closest to reproducing the observed SFRD value. It is notable, that for the other SPLs that do not overall recover the peak star formation as well, the b5 SFH implementation instead produces the best peak recovery.
The fourth panel shows the recovery of the age at which the peak in the CSFH occurs. On average the b5 SFH implementation produces the smallest offset, however the range is also very large, as the recovered SFRD is constant over around 6 Gyr with this method. The peak is frequently over- or under-estimated by up to 4 Gyr, showing the difficulty in recovering this property via SED fitting. Overall, FSPS tends to be best performing in this space.
@@ 5 Impacts of SPL ingredients
In the previous section, we highlight that the choice in SPL produces a substantial impact on the resulting properties when conducting SED fitting on a sample of galaxies. In this section, we specifically isolate the impacts that individual ingredients into the SPL have on the CSFH, by changing individual ingredients within ProGeny, as the functionality exists here to do so.
@@ 5.1 Impact of the SPL Isochrone
Figure 18: The CSFH as derived when using different isochrones within ProGeny, all assuming a Chabrier IMF and the C3K stellar atmospheres.
@@
Figure 19: The stellar mass offets resulting when using different isochrones within ProGeny, all assuming a Chabrier IMF and the C3K stellar atmospheres.
@@
The flexibility of ProGeny allows us to demonstrate the impact of purely changing the isochrone origin of the SPL. Selecting the typical isochrones used in all the studied SPLs of this work. Fig. 18 shows the derived CSFH when using the MIST (Dotter, 2016) isochrones in orange (the default in the ProGeny outputs presented in this paper), as well as the PARSEC (Marigo et al., 2017) isochrones (very similar to the Padova isochrones, also used in BC03, FSPS, and CB19), and the BaSTI isochrones (recently released, but as of yet not used in the other SPLs presented in this paper Hidalgo et al., 2018; Pietrinferni et al., 2021; Pietrinferni et al., 2024). For consistency, a Chabrier (2003) IMF is assumed for all these outputs. PARSEC produces a CSFH with a peak skewed younger than the fiducial version using MIST, whilst the BaSTI isochrones result in a substantially older CSFH. The derived spread caused by these isochrones is generally consistent with the spread between different SPLs, indicating that a large fraction of the variation caused by the choice of SPL derived from the isochrone selection. The CSFH derived from all isochrones is exactly consistent within the last 2Gyr, showing that any discrepancies from isochrones are only expressed for older populations.
@@
Fig. 19 presents the relative impact on the inferred stellar masses, highlighting that this is low on average (the maximum offset is $-0.022$ dex between PARSEC and MIST). We note though that while the median offset between BaSTI and MIST is low, is is mass-dependent, with the masses of high-mass galaxies predicted to be higher by BaSTI, but lower at lower stellar masses. In terms of the impact on SFR, the PARSEC-recovered SFRs have a median offset of 0.066 dex greater than MIST, while BaSTI and MIST produce much more consistent SFR values. On average both PARSEC and BaSTI produce younger half-mass ages than MIST (median offsets of $-0.035$ and $-0.023$ respectively). While for PARSEC this is consistent with the trend seen in the CSFH in Fig. 18, the CSFH being skewed older when using the BaSTI isochrones is caused by the mass-dependent behaviour, where for galaxies with $\log(M_{*}/{\rm M}_{\odot})>10.5$ the half-mass ages are overestimated by around 0.05 dex.
@@
As presented by Robotham & Bellstedt (in prep), the MIST isochrones contain the largest population of models in the hot post-AGB [^12] AGB - Asymptotic Giant Branch phase, while BaSTI has a lower sampling, and PARSEC almost entirely neglects this stellar evolution phase. This is a suggestion that no post-AGB coverage (PARSEC) results in an age underestimate, whereas full post-AGB coverage (MIST) recovers older ages. The consequence of an only partial coverage (BaSTI) is a mass-dependent bias of the ages, where massive galaxies are overestimated in age, and low-mass galaxies are underestimated in age. Fig. 18 hence shows the impact of including this stellar evolution phase on SED fitting very clearly.
5.2 Impact of the SPL Stellar Atmosphere
@@
Figure 20: The CSFH as derived when using different stellar atmospheres within ProGeny, all assuming a Chabrier IMF and the MIST isochrones.
@@
Figure 21: The stellar mass offets resulting when using different stellar atmospheres within ProGeny, all assuming a Chabrier IMF and the MIST isohrones. The fiducial version compared against uses the C3K stellar atmospheres.
@@
The stellar atmospheres used in any given SPL determine the selected spectrum used for any stellar property specified by the isochrones.
@@
As described in detail by Robotham & Bellstedt (in prep), ProGeny has been constructed in such a way as to combine multiple sets of astmospheres to facilitate maximal coverage of the isochrones. This is facilitated by setting a ‘base’ set of templates to cover most stars, which can be supplemented by an ‘extend’ set from a different set of stellar atmospheres to improve coverage. ProGeny supports various different publicly-available atmosphere sets, including C3K (Conroy et al., 2014), MILES (Sánchez-Blázquez et al., 2006), PHOENIX (two versions of which are available Allard et al., 2012; Husser et al., 2013), and BaSeL (Lejeune et al., 1997). MILES is supported by FSPS, and BaSeL is used broadly across differens SPLs, including BC03, BPASS, M05, CB19 and FSPS. In the fiducial implementation of ProGeny in this paper, the ‘base’ atmospheres used are from C3K, with the ‘extend‘ set selected from the Allard et al. (2012) version of the PHOENIX spectra. Using the MIST isochrones and Chabrier IMF fixed for each implementation, we vary the input ‘base’ stellar atmospheres to quantify the impact of these atmosphere sets on the derived properties from SED fitting. When modifying the atmosphere set being used, it is the ‘base’ being modified, and in all cases the Allard et al. (2012) version of the PHOENIX spectra are used as the supplemented ‘extend’ set. For our PHOENIX version, the base set is selected as the Husser et al. (2013) version of the PHOENIX spectra.
@@
As in Fig. 18, Fig. 20 shows the impact on the resulting CSFH that purely changing the atmosphere sets causes when used in SED fitting. The impact on the CSFH is similar to the impact of a constant metallicity assumption seen in the SED fitting configurations. Certainly the variation is greater than that recovered through sampling uncertainty in the SED fitting. The fiducial C3K stellar atmospheres recover the oldest CSFH, while PHOENIX, BaSeL, and MILES all recovering a CSFH that slightly overestimates the SFRD below 7 Gyr in lookback time. The present-day SFRD value recovered using C3K is slightly lower than the value recovered with all other atmosphere sets, however all still within the uncertainty of the D’Silva et al. (2023) observational measurement. Fig. 23 shows the resulting offsets in stellar masses, which are seen to be very low (the biggest discrepancy is between C3K and MILES, producing a median offset of 0.045 dex). Correspondingly, the resulting offsets in SFR are low (a maximum offset of 0.031 dex between C3K and MILES, where MILES recovers larger SFRs on average), and the maximum offset in age is between C3K and MILES, at 0.051 dex (consistent with the trend seen in the CSFH in Fig. 20).

### v3

@@ 4.5 Impact on Derived Cosmic Star Formation History
We can use the volume-limited nature of our GAMA sample to compute the CSFH using the individual SFHs in the same manner as outlined by Carnall et al. (2019); Leja et al. (2019); Bellstedt et al. (2020b); Bellstedt et al. (2024), as a manner of assessing the systematic bias of the SFHs. The CSFH is well understood from high-$z$ observations from a number of works (including Madau & Dickinson, 2014; Driver et al., 2018; D’Silva et al., 2023), and therefore any deviation in the forensic131313‘Forensic’ here refers to the process of constructing a historical property (the CSFH) from present-day galaxies. reconstructed CSFH and the observed one points to underlying age biases in the SED fitting process.
The methodology of computing the CSFH from individual SFHs is identical to that outlined by Bellstedt et al. (2024). Briefly, each individual SFH is shifted to be in lookback time units (necessary because individual SFHs are computed relative to the epoch of the galaxy), and then all SFHs are summed to produce a total SFH. Incompleteness in the sample is accounted for by scaling the SFH contributions of individual galaxies with $V_{\rm max}<V_{\rm sample}$ by a volume correction of $V_{\rm sample}/V_{\rm max}$, using the ProSpect-derived $V_{\rm max}$ values. If $V_{\rm max}\geq V_{\rm sample}$, then no correction is applied. Finally, the summed SFH is divided by the total volume of the sample to derive the star formation rate density (SFRD) with lookback time. We only present the CSFH at $z>0.06$, as below $z=0.06$ a changing number of galaxies contributes to the CSFH, resulting in numerical artefacts.
@@
The top panel of Fig. 19 presents the derived CSFH using the fiducial implementation for all six SPLs analysed in this work, as compared with the compilation of measurements that derive the star formation rate density using either UV or FIR luminosity functions by Madau & Dickinson (2014) [^14] Measurements included in this compilation come from a suite of studies: Sanders (2003); Takeuchi et al. (2003); Wyder et al. (2005); Schiminovich et al. (2005); Dahlen et al. (2007); Reddy & Steidel (2009); Robotham & Driver (2011); Magnelli et al. (2011); Magnelli et al. (2013); Cucciati et al. (2012); Bouwens et al. (2012a, b); Schenker et al. (2013); Gruppioni et al. (2013) and the more recent measurement by D’Silva et al. (2023) that uses stellar population modelling of large galaxy redshift surveys to derive the star formation rate density. The agreement between recovered CSFH curves using different SPLs is relatively good until 6 Gyr, beyond which the curves are more divergent, pointing to the underlying disagreements and uncertainties. This is consistent with the conclusions reached by Bravo et al. (2022), who determined that the recovery of galaxy properties from SED fitting was accurate to around 6 Gyr, beyond which modelling choices drove the recovered properties. It is notable that the CSFH derived using BC03, FSPS, and M05 all recover a substantial peak beyond 12 Gyr in lookback time, which is not reflected in the high-redshift of Madau & Dickinson (2014) and D’Silva et al. (2023) presented in the figure. A much smaller peak at this same epoch is produced by ProGeny, demonstrating that in four of the six SPLs studied in this work there is an over-assignment of ancient stars to galaxies. Furthermore, the observed peak of the CSFH at a lookback time of $\sim$10 Gyr is not well recovered by BC03, M05, CB19 and ProGeny, instead being much flatter between 6-12 Gyr. The overall shape of the CSFH is best recovered using FSPS, however the overall normalisation is slightly low (a consequence of the very high ancient peak).
@@
The remaining panels of Fig. 19 present the derived CSFH using all tested SPLs with the alternative implementations explored in this work. From a visual comparison between the derived CSFH curves and the underlying observational measurements, it is clear that the fiducial implementation is better capable of recovering the observed CSFH. With the change to assuming a constant Z, it can be seen that the CSFH curves tend to start skewing toward younger peaks, which is a trend that continues further when moving to the $Z_{\odot}$ implementation in the third panel. It is also notable that there is a clear artefact recovered in the $Z_{\odot}$ implementation, where the SFRD values are overestimated around a Gyr in the past. The only SPL for which this behaviour is not present is M05. While the overall recovery from the b5 SFH configuration in the bottom panel is similar to the fiducial, the impact of the stepwise parametrisation is clear in that the gradients are unable to be recovered. Recent work by Mosleh et al. (2025) studying the recovery of SFHs in mock high-z galaxies showed that stepwise parametrisations were less able to recover bursts, and tended to overestimate SFRs in the earliest parts of the SFH. This is linked to the difficulty of these SFHs to recover the gradients of the SFHs. In the b5 SFH implementation it is noteworthy that the lowest-$z$ bin of the CSFH displays much more discrepancy between the SPL values than the other SED-fitting implementations. Note that the most recent burst of star formation available with the b5 SFH implementation hardly appears in this plot, as the CSFH only focuses on the region of lookback time that is covered by all galaxies, beyond $z=0.06$. This is an indication that the relative decoupling of the SFR values in neighbouring bins acts to reduce the constraint on slightly older stellar populations. In general, the quality of CSFH recovery by the b5 mode is similar to the recovery presented by Leja et al. (2019) using a stepwise parametric SFH with the code Prospector, where it was shown that the selection of a prior can also have a large impact on the SFH recovery.
@@
A numerical comparison of the overall CSFH recovery is shown in Fig. 20. The top panel presents the mean absolute deviation (MAD) from the D’Silva et al. (2023) CSFH for each SPL/configuration combination, computed over the lookback time range for which there are observations ($<$12.5 Gyr). The ProGeny fiducial run has the overall lowest MAD (only marginally better than FSPS). In general the fiducial configuration is best capable of replicating the CSFH, as indicated by the fact that the MAD averaged over SPLs (rightmost panel of Fig. 20) is the lowest for the fiducial implementation. BPASS is the only SPL for which the fiducial implementation does not produce the preferred CSFH. The different age–metallicity behaviour of BPASS as shown in Fig. 18 and discussed in Sec 4.4.1 is deemed responsible for this, where the older ages are not preferred (likely due to the UV behaviour of high-metallicity templates) in the fiducial run when there is sufficient flexibility in metallicity.
The second panel of Fig. 20 presents the offset of the SFRD at lookback times of 0. The grey shaded region indicates the range covered by the observational uncertainty of D’Silva et al. (2023). In general, the $Z_{\odot}$ implementation can be seen to produce too much star formation, being the only configuration that over-predicts the current SFRD. However, this value is in best agreement with that of D’Silva et al. (2023). Conversely, the fiducial, constant Z, and b5 SFH implementations tend to produce underestimates. Overall, M05, BC03 and ProGeny produce a present-day SFRD that is most consistent with observations. This is perhaps unsurprising, as BC03 was used to infer the observational results by D’Silva et al. (2023).
The third panel shows the offset of the SFRD at the observed peak of the CSFH. It is notable that the magnitude of the offsets at $\sim$10 Gyr in lookback time are substantially larger than present day values (by a factor of $\sim$6), demonstrating the increased difficulty of constraining star formation at older times. All SPLs underestimate the value of the peak, however FSPS is the closest to reproducing the observed SFRD value on average. While the single implementation that gets closer to the SFRD peak is the CB19 b5 implementation, the other CB19 implementations (in particular those of the $Z_{\odot}$ and constant Z) have a much poorer reproduction of the SFRD peak. It is notable, that for the other SPLs that do not overall recover the peak star formation as well, the b5 SFH implementation instead produces the best peak recovery.
The fourth panel shows the recovery of the age at which the peak in the CSFH occurs. On average the b5 SFH implementation produces the smallest offset, however the range is also very large, as the recovered SFRD is constant over around 6 Gyr with this method. The peak is frequently over- or under-estimated by up to 4 Gyr, showing the difficulty in recovering this property via SED fitting. Overall, FSPS and BPASS tend to be best performing in this space.
@@ 5 Impacts of SPL ingredients
In the previous section, we highlight that the choice in SPL produces a substantial impact on the resulting properties when conducting SED fitting on a sample of galaxies. In this section, we specifically isolate the impacts that individual ingredients into the SPL have on the CSFH, by changing individual ingredients within ProGeny, as the functionality exists here to do so. This expands on sec. 4 of Robotham & Bellstedt (2024), which presents the impact of these individual ingredients on the ProGeny SPL and the SSPs generated. While the focus of this comparison is to assess the impact at the SED fitting stage, we highlight that a wealth of studies have been conducted to compare the outputs of simple stellar population models and the ingredients used to craft them on sources like globular clusters (for example Charlot et al., 1996; Maraston, 1998; Thomas et al., 2003; Maraston, 2005; Percival & Salaris, 2009; Martins et al., 2019; Coelho et al., 2020).
@@ 5.1 Impact of the SPL Isochrone
Figure 21: The CSFH as derived when using different isochrones within ProGeny, all assuming a Chabrier IMF and the C3K stellar spectra.
@@
Figure 22: The stellar mass offets resulting when using different isochrones within ProGeny, all assuming a Chabrier IMF and the C3K stellar spectra.
@@
The flexibility of ProGeny allows us to demonstrate the impact of purely changing the isochrone origin of the SPL. Selecting the typical isochrones used in all the studied SPLs of this work. Fig. 21 shows the derived CSFH when using the MIST (Dotter, 2016) isochrones in orange (the default in the ProGeny outputs presented in this paper), as well as the PARSEC (Marigo et al., 2017) isochrones (very similar to the Padova isochrones, also used in BC03, FSPS, and CB19, but including the COLIBRI tracks), and the BaSTI isochrones (recently released, but as of yet not used in the other SPLs presented in this paper Hidalgo et al., 2018; Pietrinferni et al., 2021; Pietrinferni et al., 2024). For consistency, a Chabrier (2003) IMF is assumed for all these outputs. PARSEC produces a CSFH with a peak skewed younger than the fiducial version using MIST, whilst the BaSTI isochrones result in a substantially older CSFH. The derived spread caused by these isochrones is generally consistent with the spread between different SPLs, indicating that a large fraction of the variation caused by the choice of SPL derived from the isochrone selection. The CSFH derived from all isochrones is exactly consistent within the last 2Gyr, showing that any discrepancies from isochrones are only expressed for older populations.
@@
Fig. 22 presents the relative impact on the inferred stellar masses, highlighting that this is low on average (the maximum offset is $-0.022$ dex between PARSEC and MIST). We note though that while the median offset between BaSTI and MIST is low, it is mass-dependent, with the masses of high-mass galaxies predicted to be higher by BaSTI, but lower at lower stellar masses. In terms of the impact on SFR, the PARSEC-recovered SFRs have a median offset of 0.066 dex greater than MIST, while BaSTI and MIST produce much more consistent SFR values. On average both PARSEC and BaSTI produce younger half-mass ages than MIST (median offsets of $-0.035$ and $-0.023$ respectively). While for PARSEC this is consistent with the trend seen in the CSFH in Fig. 21, the CSFH being skewed older when using the BaSTI isochrones is caused by the mass-dependent behaviour, where for galaxies with $\log(M_{*}/{\rm M}_{\odot})>10.5$ the half-mass ages are overestimated by around 0.05 dex.
@@
As presented by Robotham & Bellstedt (2024), the MIST isochrones contain the largest population of models in the hot post-AGB [^15] AGB - Asymptotic Giant Branch phase, while BaSTI has a lower sampling, and PARSEC almost entirely neglects this stellar evolution phase. This is a suggestion that no post-AGB coverage (PARSEC) results in an age underestimate, whereas full post-AGB coverage (MIST) recovers older ages. The consequence of an only partial coverage (BaSTI) is a mass-dependent bias of the ages, where massive galaxies are overestimated in age, and low-mass galaxies are underestimated in age. Fig. 21 hence shows the impact of including this stellar evolution phase on SED fitting very clearly.
5.2 Impact of the SPL Stellar Spectra
@@
Figure 23: The CSFH as derived when using different stellar spectra within ProGeny, all assuming a Chabrier IMF and the MIST isochrones.
@@
Figure 24: The stellar mass offets resulting when using different stellar spectra within ProGeny, all assuming a Chabrier IMF and the MIST isohrones. The fiducial version compared against uses the C3K stellar spectra.
@@
The stellar spectra used in any given SPL determine the selected spectrum used for any stellar property specified by the isochrones.
@@
As described in detail by Robotham & Bellstedt (2024), ProGeny has been constructed in such a way as to combine multiple sets of astmospheres to facilitate maximal coverage of the isochrones. This is facilitated by setting a ‘base’ set of templates to cover most stars, which can be supplemented by an ‘extend’ set from a different set of stellar spectra to improve coverage. ProGeny supports various different publicly-available stellar spectra sets, including C3K (Conroy et al., 2018), MILES (Sánchez-Blázquez et al., 2006), PHOENIX (two versions of which are available Allard et al., 2012; Husser et al., 2013), and BaSeL (Lejeune et al., 1997). A summary of the properties defining these stellar spectra sets is provided in table 3 of Robotham & Bellstedt (2024). C3K and PHOENIX are both synthetically generated (C3K derived via the ATLAS and SYNTHE codes (Kurucz, 1993), and PHOENIX is derived using the code of the same name (Hauschildt & Baron, 1999)), while MILES is empirical, and BaSeL semi-empirical. MILES is supported by FSPS, and the version used within ProGeny is taken directly from FSPS. Of particular note, is that the property gridding is adapted to be consistent with BaSeL, and furthermore the wavelength range is extended to that used in BaSeL as well (different from the wavelength extension implemented in EMILES Vazdekis et al. 2016). BaSeL is used broadly across different SPLs, including BC03, BPASS, M05, CB19 and FSPS. In the fiducial implementation of ProGeny in this paper, the ‘base’ stellar spectra used are from C3K, with the ‘extend‘ set selected from the Allard et al. (2012) version of the PHOENIX spectra. Using the MIST isochrones and Chabrier IMF fixed for each implementation, we vary the input ‘base’ stellar spectra to quantify the impact of these stellar spectra sets on the derived properties from SED fitting. When modifying the stellar spectra being used, it is the ‘base’ being modified, and in all cases the Allard et al. (2012) version of the PHOENIX spectra are used as the supplemented ‘extend’ set. For our PHOENIX version, the base set is selected as the Husser et al. (2013) version of the PHOENIX spectra.
@@
As in Fig. 21, Fig. 23 shows the impact on the resulting CSFH that purely changing the stellar spectra sets causes when used in SED fitting. The impact on the CSFH is similar to the impact of a constant metallicity assumption seen in the SED fitting configurations. Certainly the variation is greater than that recovered through sampling uncertainty in the SED fitting. The fiducial C3K stellar spectra recover the oldest CSFH, while PHOENIX, BaSeL, and MILES all recover a CSFH that slightly overestimates the SFRD below 7 Gyr in lookback time. The present-day SFRD value recovered using C3K is slightly lower than the value recovered with all other stellar spectra sets, however all still within the uncertainty of the D’Silva et al. (2023) observational measurement. Fig. 27 shows the resulting offsets in stellar masses, which are seen to be very low (the biggest discrepancy is between C3K and MILES, producing a median offset of 0.045 dex). Correspondingly, the resulting offsets in SFR are low (a maximum offset of 0.031 dex between C3K and MILES, where MILES recovers larger SFRs on average), and the maximum offset in age is between C3K and MILES, at 0.051 dex (consistent with the trend seen in the CSFH in Fig. 23).
Identifying the specific origin for these discrepancies is difficult, due to the large number of steps between spectrum selection and the final galaxy properties being derived. In each of these steps, a multitiude of small differences can have cascading effects. C3K produces a CSFH that is most unique (whereas PHOENIX, BaSeL, and MILES have CSFHs that are more similar, in particular within the last 7 Gyr), however based on the main properties of the stellar spectra sets (wavelength range, theoretical/observational origin, metallicity range etc) there is nothing that particularly distinguishes C3K.